OpenStudy (anonymous):

What are the factors of 3x2 + 10x + 8? (3x + 8)(x + 1) (3x + 1)(x + 8) (3x + 2)(x + 4) (3x + 4)(x + 2) <- my answer correct me if I'm wrong thanks

OpenStudy (whpalmer4):

Multiply it out to check: \[(3x+4)(x+2) = 3x(x+2) + 4(x+2) = 3x^2 + 6x + 4x + 8\]\[\qquad= 3x^2 + 10x + 8\checkmark\]

OpenStudy (anonymous):

thank you

OpenStudy (whpalmer4):

You're welcome!